### v2

# Do Large Language Models Mirror Cognitive Language Processing?

Abstract
Large Language Models (LLMs) have demonstrated remarkable abilities in text comprehension and logical reasoning, indicating that the text representations learned by LLMs can facilitate their language processing capabilities. In cognitive science, brain cognitive processing signals are typically utilized to study human language processing. Therefore, it is natural to ask how well the text embeddings from LLMs align with the brain cognitive processing signals, and how training strategies affect the LLM-brain alignment? In this paper, we employ Representational Similarity Analysis (RSA) to measure the alignment between 23 mainstream LLMs and fMRI signals of the brain to evaluate how effectively LLMs simulate cognitive language processing. We empirically investigate the impact of various factors (e.g., pre-training data size, model scaling, alignment training, and prompts) on such LLM-brain alignment. Experimental results indicate that pre-training data size and model scaling are positively correlated with LLM-brain similarity, [^1] For notational simplicity, we refer to the similarity between LLM representations and brain cognitive language processing signals as LLM-brain similarity. and alignment training can significantly improve LLM-brain similarity. Explicit prompts contribute to the consistency of LLMs with brain cognitive language processing, while nonsensical noisy prompts may attenuate such alignment. Additionally, the performance of a wide range of LLM evaluations (e.g., MMLU, Chatbot Arena) is highly correlated with the LLM-brain similarity.
1 Introduction
Large language models, e.g., chatGPT (OpenAI,, 2023), have demonstrated linguistic capabilities that progressively approach human-level language comprehension and generation. The text representations learned by LLMs transform text into a high-dimensional semantic space (Mikolov et al.,, 2013). In neuroscience, brain cognitive processing signals, e.g., EEG, fMRI signals, have been widely acknowledged to be able to represent a snapshot of cognitive representations of language in human brains (Xu et al.,, 2016). Such signals record human brain activities during the cognitive processing of language and play a key role in revealing the internal workings of human language cognition (Friederici,, 2011).
Figure 1: Diagram of the proposed LLM-brain similarity estimation framework. ‘S’ denotes sentence representations from different modalities. ‘$\rho$’ denotes the pearson correlation coefficient.
Thus, it is natural to ask: how “similar” are these two different modalities of language representations (LLMs representations and cognitive language processing signals in the brain) ? or do LLMs mirror human cognitive language processing? Answering these questions might facilitate us to interpret LLMs from the perspective of cognitive science.
Previous studies have explored the assessment of consistency between representations learned by pre-trained language models and cognitive processing signals (Anderson et al.,, 2013, 2015; Xu et al.,, 2016; Hollenstein et al., 2019a, ; Abnar et al.,, 2019). Along this research line, recent studies compare human neuroimaging data with representations learned 43 pre-trained language models that are smaller than 1B-sized LLMs, revealing that language models with stronger predictive abilities for the next word exhibit greater alignment with the human brain (Schrimpf et al.,, 2021). However, current LLMs have scaled up their model size by hundreds of times (e.g. LLaMa3-400B+ (Meta,, 2024)), and been strengthened by alignment training methods such as supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) (Ouyang et al.,, 2022). These factors contribute significantly to the enhanced capabilities of LLMs (Wei et al.,, 2022; Zheng et al., 2023b, ). Nevertheless, it is yet to be tested whether the latest LLMs with scales of 7B and above maintain consistency with human cognitive data. Additionally, the impact of pre-training data size, alignment training, prompts, and other training strategies on LLM-brain similarity has not been investigated.
This paper aims to assess the extent to which LLMs mirror cognitive language processing activity and subsequently analyze the impacts of different factors (e.g., pre-training data size, model scaling, alignment training) on such LLM-brain alignment. Specifically, we adopt functional magnetic resonance imaging (fMRI) signals as human cognitive language processing signals, and RSA (Kriegeskorte et al.,, 2008), widely used in neuroscience (Connolly et al.,, 2012; Diedrichsen and Kriegeskorte,, 2017), to compute similarity scores between LLM representations and human cognitive language processing signals. The framework for LLM-brain similarity computation is illustrated in Figure 1.
Our experiments comprise 23 open-source LLMs, trained with various pre-training data size, scaling and training strategies (i.e., pre-training, SFT, RLHF), to investigate the impact of these factors on LLM-brain similarity. We further analyze the impact of prompts on the human intention understanding abilities of LLMs and the consistency between LLM sentiment and human sentiment. Additionally, we investigate the relationship between the capability of LLM evaluations and LLM-brain similarity.
Our main contributions can be summarized as follows.
  • We employ human cognitive language processing signals to assess the resemblance between LLMs and human cognition. By analyzing the similarity of 23 LLMs to human cognitive language processing, we investigate the impact of various factors, such as pre-training data size, model scaling, alignment training and prompts, on the LLM-brain similarity.
  • We find that the explicit-prompts result in a higher LLM-brain similarity compared to no-prompts, indicating that explicit prompts can enhance the alignment of LLMs with human intentions.
  • Our findings indicate that LLMs exhibit high similarity to brain cognitive language processing in positive sentiment, suggesting that LLMs may encode more positive sentimental text during training.
  • We observe a high degree of consistency between the performance of a wide range of LLM evaluations (e.g., MMLU, Chatbot Arena) and LLM-brain similarity, indicating that the LLM-brain similarity holds substantial potential to evaluate the capabilities of LLMs.
2 Related Work
Brain-Inspired NLP
Human brain contains intricate neural networks and hierarchical structures, that encompass a broader cognitive complexity and generalization, providing a more comprehensive and profound perspective to understand the performance of LLMs in language cognition (Denève et al.,, 2017; Thiebaut de Schotten and Forkel,, 2022). Thus, aligning brain cognitive processing signals with text embeddings to evaluate LLMs has remained a prominent area of research. Numerous studies explore how embedding representations of words and sentences extracted from NLP models correspond to fMRI or MEG recordings (Mitchell et al.,, 2008; Huth et al.,, 2016; Pereira et al.,, 2018). Hollenstein et al., 2019b propose a framework for cognitive word embedding evaluation called CogniVal. This framework assesses the extent to which language model representations reflect semantic information in the human brain across three modalities: eye-tracking, EEG, and fMRI. Xu et al., (2016) introduce a lightweight tool named BrainBench, designed to evaluate the word semantics in distributional models, using fMRI images corresponding to 60 concrete nouns. Toneva and Wehbe, (2019) interpret the differences among 4 language models concerning layer depth, context length, and attention types using fMRI data. By leveraging insights from attention experiments, the performance of syntax-related tasks is enhanced.
NLP-inspired Neuroscience
Leveraging the superior semantic expressiveness of text embeddings from pre-trained models, a few studies utilize the computational mapping between brain activity and text embeddings to decode human brains (Mitchell et al.,, 2008; Pereira et al.,, 2018; Nishida and Nishimoto,, 2018). Sun et al., (2019) have explored the effectiveness of different types of distributed representation models in sentence-level brain decoding based on fMRI, and found that supervised structured models exhibited remarkable performance in brain decoding. Additionally, Wehbe et al., (2014) drew an analogy between the recurrent neural network language model and the working mechanisms of the brain during discourse comprehension. They utilize the predictive performance of magnetoencephalography on hidden states at different moments to reveal what information is encoded by brain regions at specific times.
3 Methodology
To measure the extent to which LLMs align with brain cognitive language processing, we employ the RSA (Kriegeskorte et al.,, 2008) method to assess the similarity in representations between human brains and LLMs when presented with the same stimulus text. Furthermore, we investigate how LLM-brain similarity varies with different prompts and analyze resemblances between LLMs and brain cognitive language processing regarding sentimental polarity using crafted prompts.
3.1 LLM-Brain Similarity Estimation
RSA is a widely-used technique for comparing cross-modal representation spaces and has found extensive applications in cognitive neuroscience (Connolly et al.,, 2012; Diedrichsen and Kriegeskorte,, 2017). In NLP, this method is utilized to calculate the consistency between two different representation spaces encoding information (Abdou et al.,, 2019; Chrupała and Alishahi,, 2019). Moreover, it can be employed to quantify the relationship between neuroscience and neural network models (King et al.,, 2019).
We hence use RSA to measure the LLM-brain similarity. Firstly, we compute the representation dissimilarity matrix (RDM) separately for brain cognitive processing signals and LLMs embeddings. For each pair of stimulus sentences, we calculate the 1 - Pearson correlation coefficient (${\rho}$) of their representations (either LLM representations or brain cognitive processing signals) as dissimilarity, following the previous works (Dwivedi and Roig,, 2019; Luo et al.,, 2022). Consequently, both the human brain (RDMh) and the LLM (RDMm) yield a diagonal symmetric dissimilarity matrix of size ${n*n}$, where ${n}$ is the number of stimulus sentences. We formulate it as follows:
${\rm RDM}=\left[\begin{array}[]{ccc}0&\cdots&1-\rho(S_{1},S_{n})\\ \vdots&\ddots&\vdots\\ 1-\rho(S_{n},S_{1})&\cdots&0\\ \end{array}\right]$ (1)
We then estimate the similarity between RDMm and RDMh using the Pearson correlation coefficient. A high correlation score implies that representations learned by the LLM are highly aligned with brain cognitive processing signals.
${\rm Sim}=\rho({\rm RDM}_{h},{\rm RDM}_{m})$ (2)
3.2 Prompts
Instruction tuning significantly enhances the in-context learning ability of LLMs by understanding task requirements well through task definitions and corresponding examples (Ouyang et al.,, 2022; Peng et al.,, 2023). We aim to verify the effect of prompts on human intentions by examining whether appended prompts increase LLM-brain similarity.
To assess the sensitivity of the LLM-brain similarity to appended prompts, we compare the effects of three conditions: no prompts (vanilla input), explicit prompts, and noisy prompts. No-prompt-appending involves directly feeding text into LLMs, allowing the LLM to automatically continue the text without any prompts. To evaluate the effect of prompt on the same task, we prepend the phrase “Please complete the following text:” to the input text as an explicit prompt. Furthermore, we introduce a noisy prompt consisting of 5 randomly sampled English words, e.g., “Harmony illuminate umbrella freedom like.” to analyze the impact of nonsensical prompts under the same length.
3.3 Sentimental Polarities
LLMs have demonstrated significant achievements in value alignment (Han,, 2023; Ganguli et al.,, 2023), with sentimental expressions emerging as a prominent feature in their outputs. Therefore, it is important to analyze the consistency of LLMs and human sentimental tendencies to assess the alignment ability of LLMs. Current studies typically utilize psychological scales to analyze the similarity of sentimental expression between LLMs and humans (Huang et al.,, 2023; Wang et al.,, 2023). In this paper, we focus on the intrinsic representations of sentimental expressions in LLMs. We delve into the similarity between the embeddings of sentimental concepts learned by LLMs and brain cognitive language processing signals under the same concepts to investigate the consistency between LLMs and humans across different sentimental polarities. It’s worth noting that brain signals, which record physiological responses accompanied by emotions, are commonly used in psychology to analyze human sentiment. (Davidson,, 2003).
Specifically, we determine the sentimental polarities (i.e., positive vs. negative) of 180 concepts in fMRI stimulus texts using the NLTK (Bird et al.,, 2009). Subsequently, we select the top 10 concepts with the highest positive sentiment and the top 10 concepts with the highest negative sentiment to serve as proxies for positive and negative sentiment, as shown in Appendix A.2. To obtain LLM representations for the selected concepts, we generate prompts using the template “The sentiment of the word <concept> is”, focusing on extracting embeddings solely for the “<concept>” token. The similarity score for positive sentiment and negative sentiment are computed by averaging the LLM-brain similarity scores across all concepts within the each polarity.
4 Experiments
We conducted experiments on 23 mainstream LLMs to investigate the alignment between LLM representations and brain cognitive processing signals, and the impact of various factors on this alignment. Additionally, we studied the consistency between LLMs and brain cognitive language processing regarding sentimental polarity.
4.1 Cognitive Language Processing Data
We employed fMRI as brain cognitive language processing signals, which maps brain activity by detecting changes associated with blood flow. The fMRI data utilized in our experiments was from Pereira (Pereira et al.,, 2018), recorded on a whole-body 3-Tesla Siemens Trio scanner with a 32-channel head coil involving 5 adult subjects.  [^2] The dataset is publicly available at https://osf.io/crwz7/. The fMRI data from subjects who participated in all experiments were utilized in this paper. The stimulus texts presented to subjects consist of 180 isolated stimuli (concepts) and 627 continuous stimuli (natural language sentences). Due to individual variations among human subjects, we calculated LLM-brain similarity separately for each individual. The overall similarity scores were derived by averaging the individual scores. More details on the fMRI data are in Appendix A.1.
4.2 Large Language Models
We used 23 LLMs from 8 LLM families.
Amber
is a 7B LLM pre-trained on 1.3 TB of tokens, with all training codes, training parameters, and system configurations open-sourced to enhance the reproducibility and scalability of LLMs (Liu et al.,, 2023). Amber’s pre-training data is divided into 360 data chunks, and 360 checkpoints are open-sourced based on these trained data chunks. Additionally, several fine-tuned versions, including Amberchat and Ambersafe, have been released. Amberchat is fine-tuned using instruction data from WizardLM (Xu et al.,, 2023), while Ambersafe conducts direct parameter optimization (DPO) based on Amberchat to align with human preferences. In this paper, we evenly select 10 checkpoints from the publicly available 360 checkpoints to investigate the impact of pre-training data size on LLM-brain similarity.
LLaMA
is a series of open-source and powerful base language models, trained on trillions of tokens from publicly available datasets. These LLMs have demonstrated state-of-the-art performance across various benchmarks (Touvron et al., 2023a, ). We selected the 7B and 13B variants of LLaMA to investigate the impact of model scaling on LLM-brain similarity.
LLaMA2
is a collection of pre-trained and fine-tuned LLMs, with model size ranging from 7B to 70B (Touvron et al., 2023b, ). Compared to LLaMA, the pre-training data size has increased by 40%, and the context length has expanded from 2k to 4k. The fine-tuned version, LLaMA2-Chat, aligns with human preferences through SFT and RLHF. Across various benchmarks and human evaluations, LLaMA2-Chat outperforms open-source chat models, demonstrating comparable performance to mainstream closed-source models such as ChatGPT (OpenAI,, 2023) and PaLM (Chowdhery et al.,, 2023). In this paper, we selected LLMs with sizes of 7B, 13B, and 70B from the LLaMA2 family.
LLaMA3
is an improved version of LLaMA2, available in both pre-trained and instruction-tuned versions (Meta,, 2024). Compared to LLaMA2, LLaMA3 has made significant advancements, including the number of trained tokens has increased from 2TB to 15TB, the context window size has expanded from 4096 to 8192, and the vocabulary has enlarged from 32,000 to 128,000. During fine-tuning stage, LLaMA3 integrates SFT, rejection sampling, proximal policy optimization (PPO), and DPO. These enhancements have led to substantial improvements in inference capabilities, code generation, and instruction following. For this paper, we employed both 8B and 70B scales for the LLaMA3 family.
Vicuna
To investigate the effect of SFT on LLMs, we employed the Vicuna family (Zheng et al., 2023a, ). Vicuna v1.3 is fine-tuned from LLaMA using a training dataset comprising 125k user-shared conversations collected from ShareGPT.com. Additionally, Vicuna v1.5 is fine-tuned from LLaMA2, with the same training dataset as Vicuna v1.3. Both versions of Vicuna are available in 7B and 13B scales.
Table 1: The LLM-brain similarity of different model scaling and different alignment training strategies with standard deviation.
Scaling LLM Training Stage LM-brain similarity
7B Amber Pre-training 0.2038 ± 0.013
Amberchat SFT 0.2263 ± 0.015
Ambersafe SFT+DPO 0.2302 ± 0.014
LLaMA-7B Pre-training 0.2103 ± 0.016
Vicuna-7B-v1.3 SFT 0.2217 ± 0.012
LLaMA2-7B Pre-training 0.2207 ± 0.016
Vicuna-7B-v1.5 SFT 0.2346 ± 0.014
LLaMA2-7B-chat SFT+RLHF 0.2410 ± 0.017
Mistral-7B Pre-training 0.2533 ± 0.013
Mistral-7B-sft-alpha SFT 0.2481 ± 0.016
Mistral-7B-sft-beta SFT 0.2573 ± 0.015
Zephyr-7B dSFT+dDPO 0.2670 ± 0.014
8B LLaMA3-8B Pre-training 0.2540 ± 0.015
LLaMA3-8B-instruct SFT+PPO+DPO 0.2697 ± 0.014
13B LLaMA-13B Pre-training 0.2221 ± 0.017
Vicuna-13B-v1.3 SFT 0.2427 ± 0.016
LLaMA2-13B Pre-training 0.2340 ± 0.013
Vicuna-13B-v1.5 SFT 0.2563 ± 0.012
LLaMA2-13B-chat SFT+RLHF 0.2694 ± 0.013
70B LLaMA2-70B Pre-training 0.2568 ± 0.015
LLaMA2-70B-chat SFT+RLHF 0.2762 ± 0.016
LLaMA3-70B Pre-training 0.2778 ± 0.015
LLaMA3-70B-instruct SFT+PPO+DPO 0.2938 ± 0.012
Mistral-7B
is a LLM with 7B parameters that outperforms the best open-source 13B model (LLaMA2) across multiple evaluation benchmarks. This model employs grouped query attention to achieve faster inference and utilizes sliding window attention to extend the processed sequence length (Jiang et al.,, 2023).
Figure 2: The LLM-brain similarity of 10 different checkpoints on Amber. ‘ckpt’ is the abbreviation for “checkpoint”.
Mistral-7B-sft
comprises SFT versions of Mistral-7B. Mistral-7B-sft-alpha is fine-tuned on the UltraChat dataset (Ding et al.,, 2023), which consists of 1.47M multi-turn dialogues generated by GPT-3.5-Turbo. Similarly, Mistral-7B-sft-beta is fine-tuned on a cleaned version of the UltraChat dataset, containing 200k samples.
Zephyr-7B
is built upon Mistral-7B-sft-beta and trained using distilled supervised fine-tuning (dSFT) and distilled direct preference optimization (dDPO) to better alignment with human intentions in interactions. This model demonstrates comparable performance to a 70B chat model aligned using human feedback (Tunstall et al.,, 2023).
The average embeddings of all tokens from the final layer of LLMs were utilized as sentence representations. All the LLMs used in this paper are available from https://huggingface.co/models.
4.3 Results
Pre-training data size
The scaling law of LLMs plays a crucial role in the performance of LLMs. (Kaplan et al.,, 2020; Hoffmann et al.,, 2022). Hoffmann et al., (2022) suggest that with a fixed scaling of LLMs, the loss gradually decreases as the amount of pre-training data increases, thereby enhancing the performance of LLMs. In this paper, we delve into how the pre-training data size affects LLM-brain similarity scores. The LLM-brain similarity scores for Amber across 10 checkpoints representing different pre-training data sizes are shown in Figure 2. It can be observed that as the size of pre-training data increases, the LLM-brain similarity scores demonstrate gradually improvement. This indirectly suggests a negative correlation between the loss of LLMs and LLM-brain similarity scores.
Table 2: The disparity in the LLM-brain similarity between the explicit-prompt-appending and no-prompt-appending strategy on different training stages. ‘$\Delta$’ denotes the increment of the explicit-prompt-appending vs. the no-prompt-appending strategy.
Scaling LLM Training Stage $\Delta$
7B Amber Pre-training 0.0121
Amberchat SFT 0.0077
Ambersafe SFT+DPO 0.0174
LLaMA-7B Pre-training 0.0088
Vicuna-7B-v1.3 SFT 0.0153
LLaMA2-7B Pre-training 0.0061
Vicuna-7B-v1.5 SFT 0.0051
LLaMA2-7B-chat SFT+RLHF 0.0086
Mistral-7B Pre-training 0.0089
Mistral-7B-sft-alpha SFT 0.0107
Mistral-7B-sft-beta SFT 0.0096
Zephyr-7B dSFT+dDPO 0.0106
8B LLaMA3-8B Pre-training 0.0014
LLaMA3-8B-chat SFT+PPO+DPO 0.0168
13B LLaMA-13B Pre-training 0.0072
Vicuna-13B-v1.3 SFT 0.0071
LLaMA2-13B Pre-training 0.0086
Vicuna-13B-v1.5 SFT 0.0075
LLaMA2-13B-chat SFT+RLHF 0.0091
70B LLaMA2-70B Pre-training 0.0108
LLaMA2-70B-chat SFT+RLHF 0.0204
LLaMA3-70B Pre-training 0.0241
LLaMA3-70B-chat SFT+PPO+DPO 0.0351
Figure 3: The LLM-brain similarity of different prompt appending strategies on LLMs.
Scaling of LLMs
The LLM-brain similarity scores across different sizes of LLMs are illustrated in Table 1. We observe that the LLM-brain similarity increases with the size of LLMs. For example, when LLaMA2 is scaled from 7B to 70B, the corresponding LLM-brain similarity increases from 0.2206 to 0.2568. This suggests that larger LLMs better mirror cognitive language processing of the brain than smaller LLMs.
Alignment of LLMs
Alignment training can significantly enhance the alignment capability of LLMs with human values (Ouyang et al.,, 2022). Therefore, we want to explore whether LLM alignment brings changes in the representation space of LLMs to narrow the gap with representations of the brain cognitive language processing. The LLM-brain similarity results for different training stages are presented in Table 1. From these results, we observe that:
  • LLMs with alignment training exhibit higher similarity to brain cognitive processing signals compared to the pre-trained versions, with LLaMA3-70B-instruct achieving the highest performance. This indicates that alignment training guides LLMs to better align with human brain.
    Figure 4: The LLM-brain similarity of LLMs across different sentimental polarities.
  • Some LLMs, after alignment training, approach or surpass larger pre-trained LLMs. For example, the LLaMA3-8B-instruct exceeds the LLaMA-13B by 0.019, and the Zephyr-7B outperforms the LLaMA2-70B by 0.0106. This suggests that alignment training has a more significant impact on the LLM-brain similarity than model scaling.
  • Mistral-7B-sft-alpha (M-alpha) and Mistral-7B-sft-beta (M-beta) are both derived from Mistral-7B through SFT. Notably, M-alpha demonstrates lower LLM-brain similarity than Mistral-7B (0.2481 vs. 0.2573). While both models utilize SFT data from the same dataset, M-beta employs a reduced dataset by 7 fold, containing only high-quality, filtered samples. This indicates that for aligning with humans, the quality of SFT data is more important than its quantity, resonating with conclusions drawn in prior research (Touvron et al., 2023b, ).
Sensitivity of Prompts
The LLM-brain similarity results under the three different prompt appending strategies are illustrated in Figure 3. It is evident that explicit prompt appending generally outperforms the other two types of prompt appending strategies. While the difference between no-prompt-appending and noisy-prompt-appending is relatively small, no-prompt-appending tends to outperform noisy-prompt-appending for most LLMs. This suggests that explicit prompts contribute to the consistency of LLMs with brain cognitive language processing, while nonsensical noisy prompt may attenuate such alignment.
Simultaneously, across these three types of appended prompts, LLMs consistently exhibit a positive correlation between scaling and LLM-brain similarity, indicating that the relationship between scaling and LLM-brain similarity remains insensitive to appended prompts. Furthermore, we explored the sensitivity of different training stages to appended prompt. We calculated the difference between the explicit-prompt-appending and no-prompt-appending, as shown in Table 2. It is observed that, the alignment version of LLMs demonstrates a more significant improvement in LLM-brain similarity after introducing explicit prompts compared to the pre-trained version. This might be because prompt helps LLMs align more closely with human intentions, thereby enhancing LLMs’s sensitivity to prompts.
5 Discussion
5.1 Consistency with Human Sentiment
The pre-training data for LLMs are predominantly from the Internet, a realm abundant with a plethora of subjective texts infused with sentimental nuances. Thus, LLMs may inherently contain sentiment-yielding capabilities, and it is worth exploring whether the sentiment of LLMs is consistent with human moral sentiment. To probe the disparities in sentimental expressions between LLMs and humans, we conducted an analysis of the LLM-brain similarity across various sentimental polarities. The experimental results are illustrated in Figure 4. We observe that the LLM-brain similarity in positive sentiment is significantly higher than that in negative sentiment. This may be due to the data governance of training data in LLMs (Touvron et al., 2023b, ; Liu et al.,, 2023), such as toxicity filtering (Friedl,, 2023; Gargee et al.,, 2022), which retains high-quality and positive sentiment text, resulting in a deeper encoding of positive sentiment compared to negative sentiment by the LLMs.
(a) MMLU
(b) HellaSwag
(c) Chatbot Arena ELO Rating
(d) AlpacaEval
Figure 5: Correlation between the performance of LLMs on evaluations and the LLM-brain similarity.
Furthermore, alignment training enhances LLM-brain similarity in both positive and negative sentiments, indicating that LLMs encode human sentimental tendency when aligning with human sense of worth.
5.2 Correlation with LLM Evaluations
We conducted a comparative analysis between LLM-brain similarity and performance on various LLM evaluations. To assess the knowledge and capability, we selected two benchmarks: Massive Multilingual Language Understanding (MMLU) (Hendrycks et al.,, 2021) and HellaSwag (Zellers et al.,, 2019). MMLU spans 57 subjects, including STEM, humanities, and social sciences, designed to evaluate the knowledge reservoir of LLMs. HellaSwag requires LLMs to complete sentences, serving as an evaluation of commonsense reasoning abilities. For alignment evaluation, we employed Chatbot Arena Elo Rating (Zheng et al., 2023a, ), which uses crowdsourced methods for adversarial assessments of LLM outputs, and AlpacaEval 2.0 (Li et al.,, 2023), which adopts powerful LLMs such as GPT-4 to evaluate LLMs. All performance results of these LLM evaluations are from publicly available benchmarks.
Figures 5(a) and 5(b) respectively illustrate the correlation between LLM-brain similarity and the two knowledge and capability evaluations. The results suggest a positive correlation between LLM-brain similarity and performance on these evaluations, aligning with similar findings from earlier studies (Hollenstein et al., 2019b, ). An intriguing finding is that LLM-brain similarity not only assesses the consistency between LLMs and human brain but also evaluates the capabilities of LLMs.
The correlation between LLM-brain similarity and alignment evaluations is illustrated in Figures 5(c) and 5(d). The results reveal that, with the exception of LLaMA3-70B-instruct, the LLM-brain similarity and alignment capability of LLMs generally exhibit a positive correlation. LLaMA3-70B-instruct consistently achieves the highest performance in both LLM-brain similarity and alignment evaluations, significantly surpassing the linear trend line in the figures. This suggests that, although LLaMA3-70B-instruct excels in addressing artificially constructed alignment evaluations, its intrinsic representation still maintains a perceptible distance from the human brain.
6 Conclusions
This paper has presented a framework to estimate how well LLMs mirror cognitive language processing of the brain, quantified by the LLM-brain similarity. We have investigated the impact of pre-training data size, model scaling, alignment training, and prompts on the LLM-brain similarity, and explored the consistency of LLMs with humans in sentimental polarities. Experimental results reveal that expanding the size of pre-training data, scaling up models, and employing alignment training contribute to enhancing LLM-brain similarity, emphasizing the importance of high-quality SFT data in elevating LLM-brain similarity. Moreover, explicit prompts aids LLMs in understanding human intentions, and alignment training enhances the sensitivity to prompts. Notably, LLMs exhibit a stronger resemblance to humans in positive sentiment. The strong correlation between LLM-brain similarity and various LLM evaluations suggests that the proposed LLM-brain similarity possesses generalization and rationality, which could serve as a new way to evaluate LLMs from the cognitive perspective.
Appendix A Appendix
A.1 Cognitive Language Processing Data
The temporal resolution of fMRI is two seconds, meaning that under continuous stimuli, each scan covers multiple words (Pereira et al.,, 2018). The fMRI data consist of neural activity representations in cubic millimeter-sized voxels, with each fMRI recording containing a substantial number of voxels. We flattened the 3-dimensional fMRI images into 1-dimensional vectors and randomly selected 1,000 voxels as the brain cognitive language processing signal for each subject.
A.2 Sentimental Concepts
Table 3: 10 positive concepts vs. 10 negative concepts, serving as proxies to positive and negative sentiment.
      Positive       Negative
      great       war
      charming       sin
      successful       stupid
      pleasure       prison
      laugh       pain
      elegance       liar
      kindness       angry
      smiling       damage
      accomplished       sad
      impress       poor
A.3 Limitations
This study, relying on text representations from open-source LLMs, is unable to assess the LM-brain similarity of closed-source LLMs such as ChatGPT. We aim for future investigations to extend these findings to a broader spectrum of LLMs. Additionally, the fMRI stimulus texts used in this paper are exclusively in English, limiting the generalization of the observed relationships between LLMs and human cognition to other linguistic environments. Future endeavors should explore LM-brain similarity in LLMs across diverse languages.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction sections of the paper accurately reflect the contributions of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of this paper are presented in Appendix A.3.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The section 4.3 of this paper provides comprehensive evidence for all the conclusions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The section 3 of this paper provides all the information necessary for experimental replication.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The data and LLMs used in this paper are publicly available, and links are provided in the section 4.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experimental details can be found in the section 4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The standard deviation of the experimental results is provided in section 4.3. We can observe minimal variance among individuals, which to some extent indicates the reliability of the experimental results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Extracting embeddings from LLMs in this paper requires consideration of computer resource usage. The provided link in subsection 4.2 mentions the required resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The human brain cognitive processing data employed in this paper are publicly accessible and devoid of any personal information of subjects. The purpose of utilizing these signals is to investigate the extent to which LLMs align with human brain cognitive language processing. All LLMs used in the experiments are also publicly accessible.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper investigates the correlation between LLMs and human brain cognitive processing signals, without delving into societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All data and models utilized in this paper are publicly available and appropriately cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will release the code publicly after the paper is accepted.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The data we used are all publicly available and do not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.